A chemist has one solution that is 40 percent pure glycoside and another solution that is 80 percent pure glycoside. How many pounds of each solution must be used to produce 100 pounds of a solution that is 60 percent glycoside?
Answer by checkley71(8403)

.4X+.8(100-X)=.6*100
.4X+80-.8X=60
-.4X=60-80
-.4X=-20
X=-20/-.4
X=50 POUNDS OF 40% ARE NEEDED.
100-50=50 POUNDS OF 80% IS NEEDED.
PROOF
.4*50+.8*50=.6*100
20+40=60
60=60